OpenStudy (anonymous):

second derivative of 3x^2 + y^2 = 2 using implicit differentiation

OpenStudy (anonymous):

well isnt the first derivative 6x+2yy'=0

OpenStudy (perl):

6x + 2y*y' = 0 , correct

OpenStudy (perl):

y' = -6x / 2y y' = -3x/y

OpenStudy (perl):

now take the derivative of y ' , to get the second derivative

OpenStudy (gorv):

itzz just the simple derivative